Circular No. 8377

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2004dj IN NGC 2403 S. Nakano, Sumoto, Japan, reports the discovery by K. Itagaki, Teppo-

cho, Yamagata, Japan, of a bright supernova (mag 11.2, roughly V) on at least ten unfiltered CCD frames taken around July 31.76 UT using a 0.60-m f/7 reflector. The new object is located at α = 7^{h}37^{m}17^{s}.02, δ = +65^{o}35′57′′.8 (equinox 2000.0), which is roughly 160′′ east and 10′′ north of the nucleus of NGC 2403 (the nucleus being difficult to define on his images). Additional CCD magnitudes of SN 2004dj by Itagaki: 2002 Sept. 19, [18.5; Oct. 11, [18.5; 2004 Aug. 1.45, 11.3. Visual magnitude estimates: Aug. 1.910, 11.5 (R. J. Bouma, Groningen, The Netherlands); 2.031, 11.5 (M. Lehký, Hradec Králové, Czech Republic); 2.062, 11.5 (K. Hornoch, Lelekovice, Czech Republic); 2.910, 11.6 (Bouma).

SUPERNOVA 2004dk IN NGC 6118 Further to IAUC 8375, J. Graham and W. Li report the LOSS discov-

ery, on an unfiltered KAIT image taken on Aug. 1.19 UT, of an apparent supernova (mag ∼ 17.6) located at α = 16^{h}21^{m}48^{s}.93, δ = −2^{o}16′17′′.3 (equinox 2000.0), which is 4′′.4 east and 43′′.5 north of the nucleus of NGC 6118. Additional approximate KAIT magnitudes for SN 2004dk: July 25.20, [19.5; 27.19, [18.5; 29.21, [18.0; 30.19, 18.6; 31.20, 18.0.

SUPERNOVAE 2004dl AND 2004dm M. Riello, European Southern Observatory (ESO), reports (together

with G. Pignata, M. T. Botticella, F. Patat, S. Benetti, E. Cappellaro, M. Turatto, and S. Valenti) the discovery of two supernovae on CCD frames taken during a search performed with the Max-Planck-Institut 2.2-m telescope (+ Wide Field Imager) at the ESO.

SN  2004 UT  α_{2000}  δ_{2000}  R  z Type
2004dl July 14 14^{h}11^{m}11.^{s}50 −12^{o}02′36′′.9 21.2 0.250 Ia
2004dm July 14 22 43 21.36 −40 19 46.7 21.6 0.225 Ib

Nothing was visible at the above positions on reference images taken with the same equipment in May 2004 and November 2001 (limiting mag R ∼ 23.5). Spectroscopic observations (range 600–1100 nm), obtained on July 22 with the ESO 8.2-m ANTU telescope (+ FORS2), show that, from a comparison with SNe 1992A and 2000H, the new objects appear both to be ∼ 2 weeks past maximum light.

2004 August 2  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green